Use the following information to answer questions 17-20.

A large hospital uses a certain intravenous solution that it maintains in inventory. Assume the hospital uses reorder point method to control the inventory of this item. Pertinent data about this item are as follows:

------------------------------------------------------------

Forecast of demanda = 1,000 units per week

Forecast errora, std. dev. =100 units per week

Lead time = 4 weeks

Carrying cost = 25 % per year

Purchase price, delivered = $52 per unit

Replenishment order cost = $20 per order

Stockout cost = $10 per unit

In-stock Probability during the lead time =90%

a Normally distributed

------------------------------------------------------------

Due to possible rounding effect, please pick the closest number in the following options.

Question 17 (1 point)

Question 17 Saved

What is the reorder point?

Question 18 (1 point)

Question 18 Saved

If the hospital orders 400 units each time, what’s the expected number of units out of stock annually?

Question 18 options:

1248

2912

500

1000

Save

Use the following information to answer questions 17-20.

A large hospital uses a certain intravenous solution that it maintains in inventory. Assume the hospital uses reorder point method to control the inventory of this item. Pertinent data about this item are as follows:

------------------------------------------------------------

Forecast of demanda = 1,000 units per week

Forecast errora, std. dev. =100 units per week

Lead time = 4 weeks

Carrying cost = 25 % per year

Purchase price, delivered = $52 per unit

Replenishment order cost = $20 per order

Stockout cost = $10 per unit

In-stock Probability during the lead time =90%

a Normally distributed

------------------------------------------------------------

Due to possible rounding effect, please pick the closest number in the following options.

Question 19 (1 point)

Question 19 Saved

If the hospital orders 400 units each time, what’s the total annual costs (holding cost + ordering cost + stock-out cost) excluding purchasing costs?

Question 19 options:

10000

21008

31008

42016

Save

Use the following information to answer questions 17-20.

A large hospital uses a certain intravenous solution that it maintains in inventory. Assume the hospital uses reorder point method to control the inventory of this item. Pertinent data about this item are as follows:

------------------------------------------------------------

Forecast of demanda = 1,000 units per week

Forecast errora, std. dev. =100 units per week

Lead time = 4 weeks

Carrying cost = 25 % per year

Purchase price, delivered = $52 per unit

Replenishment order cost = $20 per order

Stockout cost = $10 per unit

In-stock Probability during the lead time =90%

a Normally distributed

------------------------------------------------------------

Due to possible rounding effect, please pick the closest number in the following options.

Question 20 (1 point)

Question 20 Saved

If the lead time is normally distributed with a mean of 4 weeks and a standard deviation of 0.5 weeks, what’s the reorder point?

Question 20 options:

4689

4129

5188

6000

Save

Previous PageNext Page
Answer #1

Reorder Point = Mean Demand*Lead Time + z*SD of Demand*(Lead Time)^(1/2)

ROP = 1000*4 + 1.281552*100*2

ROP = 4256.31

ROP = 4256

_______________________________________________________________________

Normal loss function, L(Z) corresponding to 90% service level = 0.048

Expected shortage per cycle (ESC) = L(Z) * σ * (L)^(1/2) = 0.0473*100*(4)^(1/2) = 9.6

Annual shortage = ESC x no. of cycles annually = 9.6 x (52000/400) = 1248

__________________________________________________________________________

Shortage cost = Annual shortage x shortage cost per unit = 1248*10 = 12480

Order quantity, Q = 400
Holding cost per unit per annum, H = 52 * 25% = 13
Annual demand, D = 1000 x 52 = 52,000
Ordering cost, S = 20

Cycle stock holding cost = Q * H/ 2 = 400*13/2 = 2600

Ordering cost = (D/Q) * S = (52000 / 400) * 20 = 2600

Safety stock cost = Safey stock * H = 1.28*100*SQRT(4) * 13 = 3328

Total cost = $2,600 + $2,600 + $12,480 + $3,328 = $21,008

__________________________________________________________________________

L = 4
σL = 0.5

The following formula for ROP will apply.

ROP = d * L + Z * √(σ2.L + d2L2)= 1000*4 + 1.28*SQRT((100^2)*4 + (1000*0.5)^2) =4,689